# Stealthy Patch-Wise Backdoor Attack in 3D Point Cloud via Curvature Awareness

### v2

@@ Abstract
Backdoor attacks pose a severe threat to deep neural networks (DNNs) by implanting hidden backdoors that can be activated with predefined triggers to manipulate model behaviors maliciously. Existing 3D point cloud backdoor attacks primarily rely on sample-wise global modifications, which suffer from low imperceptibility. Although optimization can improve stealthiness, optimizing sample-wise triggers significantly increases computational cost. To address these limitations, we propose the Stealthy Patch-Wise Backdoor Attack (SPBA), the first patch-wise backdoor attack framework for 3D point clouds. Specifically, SPBA decomposes point clouds into local patches and employs a curvature-based imperceptibility score to guide trigger injection into visually less sensitive patches. By optimizing a unified patch-wise trigger that perturbs spectral features of selected patches, SPBA significantly enhances optimization efficiency while maintaining high stealthiness. Extensive experiments on ModelNet40 and ShapeNetPart further demonstrate that SPBA surpasses prior state-of-the-art backdoor attacks in both attack effectiveness and resistance to defense methods.
@@ 1 Introduction
Point clouds serve as a highly effective data structure for representing 3D geometric data [6], finding increasing utility in safety-critical applications including autonomous driving [11], healthcare [15, 32], and robotics [38]. With this expanding deployment, the security of point cloud models becomes paramount, particularly concerning their vulnerability to backdoor attacks [10]. In these attacks, adversaries inject a pre-defined trigger into a small portion of training data. This manipulation causes the backdoored model, upon training, to consistently misclassify samples containing the trigger into a target class, while maintaining accurate predictions for benign inputs. A well-designed trigger remains inconspicuous within the data yet potently activates malicious behavior in the backdoored model. Therefore, trigger design is crucial for both the effectiveness and stealth of backdoor attacks [24, 3, 12, 28].
@@
To this end, we aim to design a novel and highly imperceptible backdoor attack trigger that solely leverages local context information for subtle local modifications. In the 2D image domain, local triggers are typically designed as unified patch-wise triggers [5, 21], which blend seamlessly into diverse image backgrounds. However, the absence of an inherent grid structure in 3D point clouds makes it challenging to design and inject consistent patch-wise deformation triggers without introducing noticeable anomalies or compromising geometric integrity. To address this challenge, we design the patch-wise trigger in the spectral domain rather than the raw spatial domain. Specifically, the Graph Fourier Transform (GFT) is employed to map point coordinates of local patches into the spectral domain, where the overall local geometric structure is maintained while the point-level fine-grained details are perturbed. As noted by [16], human perception varies across different regions of a 3D point cloud, with perturbations in geometrically complex areas being less perceptible. Building on this insight, we selectively filter out smooth patches and apply the patch-wise trigger exclusively to geometrically complex regions, thereby enhancing stealthiness, as illustrated in Fig. 1. Notably, the patch-wise spectral trigger requires significantly lower computational cost compared to the sample-wise spectral trigger in the previous IBAPC attack [2]. This greatly improves the efficiency and deployability of our attack method. What’s more, the enriched optimization landscape enabled by patch-wise decomposition, together with targeted patch optimization, is empirically found to be beneficial for improving attack effectiveness.
@@ 2 Related Work
Recent studies have explored backdoor attacks in 3D point clouds, proposing various trigger designs based on their unique properties [26, 10]. PointBA [10] introduced two attack strategies: PointBA-I, which injects an additional cluster of points as a trigger, and PointBA-O, which exploits a specific rotation angle as the backdoor trigger. Similar to PointBA-I, PCBA [29] embeds a spherical point cloud at an optimized location to manipulate model predictions. To further enhance stealthiness, IRBA [4] achieved sample-specific triggers by utilizing a nonlinear and local transformation. Additionally, iBA [1] generated poisoned samples using a folding-based autoencoder (AE), introducing a sample-wise trigger that consists of geometric perturbations during reconstruction. However, the above attack methods all rely on sample-wise trigger injection to generate poisoned samples, which inevitably introduces noticeable global geometric deformations or noise. To enhance trigger stealthiness, the prior spectral-domain attack method [2] optimized a sample-wise spectral trigger, which caused substantial computational cost and consequently limited both algorithmic efficiency and real-world applicability.
In contrast, our approach is the first to optimize a local patch-wise trigger for 3D point cloud backdoor attacks, achieving favorable results in terms of attack effectiveness, stealthiness, and computational efficiency. We observe that similar techniques have been employed in recent adversarial attacks [13, 30, 7, 8, 14]. However, our approach differs from these methods in two key aspects: (1) Attack paradigm: As test-time attacks, these adversarial attacks require sample-specific perturbations, which are optimized uniquely for each input in the spatial domain using region-specific optimization. In contrast, our backdoor attack is a train-time attack, requiring a unified patch-wise trigger that can be consistently embedded into local patches across different samples. This need for a local and transferable trigger motivates our spectral-domain design, ensuring that perturbations remain structurally coherent while maintaining high attack effectiveness; (2) Region selection strategy: Prior adversarial attacks leverage sample-level spectral analysis to filter and select different regions for perturbation. In contrast, our method introduces a curvature-based patch imperceptibility score and selects points based on the individual patch level.
3 Method
@@ 3.2 Patchification and Curvature-Aware Selection
Due to the unordered and irregular nature of point clouds, it is infeasible to directly segment patches for trigger injection in the same manner as in 2D images. To address this challenge, we draw inspiration from previous works [37, 34] and employ Farthest Point Sampling (FPS) and K-Nearest Neighbors (KNN) algorithms to systematically decompose each point cloud into local patches. This strategy enables a structured yet flexible partitioning of the point cloud into local patches. Each patch retains sufficient fine-grained geometric detail, supporting curvature-aware evaluation essential for spectral trigger embedding. Given an input point cloud sample $X$, FPS is first applied to select $g$ center points $X^{ct}=\{x^{ct}_{1},x^{ct}_{1},...,x^{ct}_{g}\}$, ensuring a well-distributed set of key reference points. For each center point $x^{ct}\in X^{ct}$, a local patch will be constructed by the KNN algorithm containing $k_{g}$ neighbors: $S_{i}=\mathit{K\!N\!N}(X,x_{i}^{ct})$ where $S_{i}\in\mathbb{R}^{k_{g}\times 3}$. The resulting set of patches is represented as $\{S_{1},S_{2},...,S_{g}\}$.
@@
where $\mathcal{N}_{x_{j}}$ is defined as the set of $k_{c}$ nearest neighbors of point $x_{j}$ and $\mathbf{n_{x_{j}}}$ denotes its normal vector. The imperceptibility score (IS) of each point is further computed as the standard deviation of curvature between the point $x_{j}$ and its neighboring points $\mathcal{N}_{x_{j}}$, capturing local curvature fluctuations:
@@ 3.3 Patch-wise Spectral Trigger Injection
To overcome these limitations, we propose designing local patch-wise triggers in the spectral domain [8, 14]. Specially, a selected point cloud patch $S^{\prime}\in\{S^{\prime}_{1},S^{\prime}_{2},...,S^{\prime}_{m}\}$ is first transformed into an unweighted KNN graph $G=(V,A)$. Here $V\in\mathbb{R}^{k_{g}\times 3}$ represents the set of vertices, where each vertex corresponds to a point in the patch $S^{\prime}$. The adjacency matrix $A\in\mathbb{R}^{k_{g}\times k_{g}}$ is constructed by connecting each point to its $k_{p}$-nearest neighbors, where $k_{p}$ is a hyperparameter for graph construction ($A_{ij}=1$ if point $i$ and $j$ are connected otherwise $A_{ij}=0$).
@@
$\hat{S}=\phi_{GFT}(S^{\prime})=U^{T}S^{\prime},$ (4)
where $\hat{S}\in\mathbb{R}^{k_{g}\times 3}$. Correspondingly, the Inverse GFT (IGFT) could reconstruct the spatial feature from the spectral domain:
@@ 3.4 Optimization Strategy
The backdoor learning process follows an alternating optimization scheme, where the model parameters $\theta$ and the spectral trigger $\xi$ are updated iteratively, with one being optimized while the other remains frozen. Thus, each optimization step consists of two forward and backward propagations.
@@
4 Experiments and Results
4.1 Experiment Details
Attacks Setup. We compare our proposed SPBA method against four representative point cloud attack methods: PointBA-I [10], PointBA-O [10], IRBA [4], and IBAPC [2]. To reproduce PointBA-I, we implant a spherical cluster comprising 3% of the total points, with a fixed radius of 0.05, positioned at $(0.5,0.5,0.5)$ in the coordinate space. For PointBA-O, we apply a $10^{\circ}$ rotation along the Z-axis. For IRBA and IBAPC, we follow the default configurations specified in their original papers. For our proposed SPBA, $g$ and $k_{g}$ are both set as 32 for patch decomposition, while the values of $k_{c}$ and $k_{p}$ are both set as 10 for the KNN graph construction. And $m=16$ patches are selected for trigger injection. For the regularizer function in Eq. (10), we set $\lambda_{1}=1$, $\lambda_{2}=5$ and $\lambda_{3}=1$. However, for PointNet, we adjust $\lambda_{1}$ to 0.1 to improve stability. For all attack methods, we set the poisoning rate $\rho=0.1$. The target class is chosen as Chair ($Y_{t}=8$) for ModelNe40 and Lamp ($Y_{t}=8$) for ShapeNetPart. Two Adam optimizers are employed for backdoored models training and spectral trigger optimization, with learning rates of 0.001 and 0.01, respectively. The training epoch is set to 100. All experiments were conducted on a machine equipped with an NVIDIA RTX 4090 GPU.
Evaluation Metrics. The attack performance can be evaluated by Benign Accuracy (BA) and Attack Success Rate (ASR). BA is the accuracy of benign test samples that are correctly classified. ASR represents the proportion of poisoned samples that are successfully misclassified into the target label. To quantify the imperceptibility of poisoned samples, we compute the Chamfer Distance (CD) between the poisoned and original samples. A successful backdoor attack should achieve a high ASR and lower CD while preserving BA as much as possible. For clarity, both BA and ASR are presented as percentages, and CD values in tables are scaled by a factor of 1,000.
4.2 Attack Effectiveness
The results presented in Tab. 1 provide a comparative evaluation of various backdoor attack methods on ModelNet40. As shown in Tab. 1, our proposed SPBA achieves the lowest average CD value of 1.08 among all attacks, demonstrating its state-of-the-art stealthiness. What’s more, SPBA outperforms most backdoor attacks according to BA and ASR. In particular, compared to the sample-wise spectral attack IBAPC, SPBA achieves notable improvements of 0.47% in average BA and 2.78% in average ASR, highlighting the effectiveness of our patch-wise spectral trigger. Although PointBA-I achieves the highest average ASR of 98.43%, it comes at the cost of a significantly higher average CD of 4.08, making it much more detectable. Subsequent defense experiments further confirm that the trigger used in PointBA-I is especially susceptible to detection or removal by existing defense methods. Tab. 2 shows more experimental results on the ShapeNetpart dataset. SPBA achieves a similar performance, offering the best balance between attack effectiveness and stealthiness. These results further validate the superiority of SPBA as a more effective and imperceptible backdoor attack method compared to existing attacks.
Figure 3: Visual comparison of poisoned samples from different backdoor attack methods. Our proposed SPBA preserves structural integrity and ensures smooth regions (e.g., vase) remain undisturbed, enhancing imperceptibility.

### v3

@@ Abstract
Backdoor attacks pose a severe threat to deep neural networks (DNNs) by implanting hidden backdoors that can be activated with predefined triggers to manipulate model behaviors maliciously. Existing 3D point cloud backdoor attacks primarily rely on sample-wise global modifications, which suffer from low imperceptibility. Although optimization can improve stealthiness, optimizing sample-wise triggers significantly increases computational cost. To address these limitations, we propose the Stealthy Patch-Wise Backdoor Attack (SPBA), the first patch-wise backdoor attack framework for 3D point clouds. Specifically, SPBA decomposes point clouds into local patches and employs a curvature-based imperceptibility score to guide trigger injection into visually less sensitive patches. By optimizing a unified patch-wise trigger that perturbs spectral features of selected patches, SPBA significantly enhances optimization efficiency while maintaining high stealthiness. Extensive experiments on ModelNet40 and ShapeNetPart further demonstrate that SPBA surpasses prior state-of-the-art backdoor attacks in both attack effectiveness and resistance to defense methods. The code is available at https://github.com/HazardFY/SPBA.
@@ 1 Introduction
Point clouds serve as a highly effective data structure for representing 3D geometric data [11], finding increasing utility in safety-critical applications including autonomous driving [17], healthcare [42], and robotics [49]. With this expanding deployment, the security of point cloud models becomes paramount, particularly concerning their vulnerability to backdoor attacks [16]. In these attacks, adversaries inject a pre-defined trigger into a small portion of training data. This manipulation causes the backdoored model, upon training, to consistently misclassify samples containing the trigger into a target class, while maintaining accurate predictions for benign inputs. A well-designed trigger remains inconspicuous within the data yet potently activates malicious behavior in the backdoored model. Therefore, trigger design is crucial for both the effectiveness and stealth of backdoor attacks [33, 45].
@@
To this end, we aim to design a novel and highly imperceptible backdoor attack trigger that solely leverages local context information for subtle local modifications. In the 2D image domain, local triggers are typically designed as unified patch-wise triggers [10], which blend seamlessly into diverse image backgrounds. However, the absence of an inherent grid structure in 3D point clouds makes it challenging to design and inject consistent patch-wise deformation triggers without introducing noticeable anomalies or compromising geometric integrity. To address this challenge, we design the patch-wise trigger in the spectral domain rather than the raw spatial domain. Specifically, the Graph Fourier Transform (GFT) is employed to map point coordinates of local patches into the spectral domain, where the overall local geometric structure is maintained while the point-level fine-grained details are perturbed. As noted by [23], human perception varies across different regions of a 3D point cloud, with perturbations in geometrically complex areas being less perceptible. Building on this insight, we selectively filter out smooth patches and apply the patch-wise trigger exclusively to geometrically complex regions, thereby enhancing stealthiness, as illustrated in Fig. 1. Notably, the patch-wise spectral trigger requires significantly lower computational cost compared to the sample-wise spectral trigger in the previous IBAPC attack [5]. This greatly improves the efficiency and deployability of our attack method. What’s more, the enriched optimization landscape enabled by patch-wise decomposition, together with selected patch optimization, is empirically found to be beneficial for improving attack effectiveness.
@@ 2 Related Work
2.1 Patch-wise Backdoor Attacks in 2D Images
Different from adversarial attacks [4, 46], backdoor attacks pose a significant security threat in the training process by embedding hidden triggers during the training phase, enabling adversaries to manipulate model predictions during inference [20, 8]. The effectiveness of a backdoor attack heavily depends on trigger design, which can be broadly categorized into global triggers and local patch-wise triggers. Prior works have investigated global triggers from various perspectives, including image steganography [18, 32], frequency-domain injection [6], and optimization-based approaches [40], aiming to balance stealthiness and attack effectiveness. As for local patch-wise triggers, BadNets [10] first demonstrated the feasibility of backdoor attacks in the 2D image domain with a fix patch at the corner. Since then, patch-wise triggers have become a fundamental paradigm in backdoor attacks due to their high attack success rates and stealthiness. Chen  et al. [3] introduced a blending strategy, embedding triggers with partial transparency to make them less detectable. Yuan  et al. [43] found that Vision Transformers (ViTs) are particularly vulnerable to patch-wise triggers, due to their global dependencies within patches. Some works further explored the adaptability of patch-wise trigger to various training paradigms, including self-supervised learning [29] and multimodal contrastive learning [1]. Gao  et al. [9] further improved patch-wise attacks by designing a sparse and invisible trigger through bi-level optimization, minimizing visual detectability while preserving attack effectiveness. What’s more, Li  et al. [19] adopted patch-wise triggers as a harmless and stealthy way to protect the dataset copyright.
Intuitively, perturbing only a small subset of points within a 3D object is both more stealthy and more practical, while still preserving the object’s global geometric structure. These properties make patch-wise triggers a particularly attractive option in point cloud scenarios. Nonetheless, most existing patch-wise trigger designs were developed for 2D images, and their direct adaptation to 3D point clouds is non-trivial due to the unordered and irregular nature of point cloud data. This motivates our work, which explores the design of stealthy and optimized patch-wise triggers for 3D point-cloud models.
2.2 Backdoor Attacks and Defenses in 3D Point Clouds
@@
Recent studies have explored backdoor attacks in 3D point clouds, proposing various trigger designs based on their unique properties. PointBA [16] introduced two attack strategies: PointBA-I, which injects an additional cluster of points as a trigger, and PointBA-O, which exploits a specific rotation angle as the backdoor trigger. Similar to PointBA-I, PCBA [37] embeds a spherical point cloud at an optimized location to manipulate model predictions. But the injected clusters are often visually conspicuous and can be effectively removed by augmentations. IRBA [7] improves stealth by generating sample-specific triggers through nonlinear local transformations, but it remains vulnerable to common data augmentations. IBAPC [5] optimizes a sample-wise spectral trigger, while iBA [2] requires training an extra folding-based autoencoder to generate geometric perturbations. Despite their different designs, both methods introduce significant computational cost that limits scalability and real-world applicability. SRBA [25] employs uniformly shifted input features as triggers, but this design alters the network’s input format, thereby limiting its generality. The above attack methods all rely on sample-wise trigger injection to generate poisoned samples, which inevitably introduces noticeable global geometric deformations or noise. In contrast, our approach is the first to optimize a local patch-wise trigger for 3D point cloud backdoor attacks, achieving favorable results in terms of attack effectiveness, stealthiness, and computational efficiency. Similar region-based perturbations have also appeared in recent adversarial attacks [39, 12]. Unlike these test-time, sample-specific methods, our approach is a train-time backdoor attack that embeds a unified patch-wise trigger transferable across samples. In addition, rather than relying on sample-level spectral analysis, we employ a curvature-based patch imperceptibility score for patch-level selection, ensuring efficiency and stealthiness.
On the defense side, conventional data augmentations are often employed as a defense strategy against backdoor attacks [7]. SOR [48] enhanced robustness by removing outliers from point clouds. Xiang et al. [38] proposed a reverse-engineering defense to detect the backdoored model. PointCRT [15] detected the poisoned samples by analyzing their abnormal corruption resistance. Lian et al. [21] introduced a novel cross-modal driven object restoration framework and mitigated the implanted triggers by reconstruction. Additionally, PointNCBW [35] explores watermarking in point clouds via a backdoor mechanism for positive purposes, which is beyond the scope of this work.
3 Preliminaries
@@ 3.1 Problem Statement
3.2 Threat Model
Consistent with prior backdoor attack methods [37, 5], we assume that adversaries operate under a white-box setting, where they have full access to the model architecture, can manipulate a subset of the training data, and control the training process. By injecting poisoned samples into the training set, adversaries are able to implant backdoors into the DNNs. At the inference stage, we assume that adversaries are constrained to black-box interactions, i.e., they can only issue queries to the backdoored models with arbitrary inputs.
4 Method
@@ 4.1 Patchification and Curvature-Aware Selection
Due to the unordered and irregular nature of point clouds, it is infeasible to directly segment patches for trigger injection in the same manner as in 2D images. To address this challenge, we draw inspiration from previous works [44] and employ Farthest Point Sampling (FPS) and K-Nearest Neighbors (KNN) algorithms to systematically decompose each point cloud into local patches. This strategy enables a structured yet flexible partitioning of the point cloud into local patches. Each patch retains sufficient fine-grained geometric detail, supporting curvature-aware evaluation essential for spectral trigger embedding. Given an input point cloud sample $X$, FPS is first applied to select $g$ center points $X^{ct}=\{x^{ct}_{1},x^{ct}_{1},...,x^{ct}_{g}\}$, ensuring a well-distributed set of key reference points. For each center point $x^{ct}\in X^{ct}$, a local patch will be constructed by the KNN algorithm containing $k_{g}$ neighbors: $S_{i}=\mathit{K\!N\!N}(X,x_{i}^{ct})$ where $S_{i}\in\mathbb{R}^{k_{g}\times 3}$. The resulting set of patches is represented as $\{S_{1},S_{2},...,S_{g}\}$.
@@
where $\mathcal{N}_{x_{j}}$ is defined as the set of $k_{c}$ nearest neighbors of point $x_{j}$ and $\mathbf{n_{x_{j}}}$ denotes its normal vector. The imperceptibility score (IS) of each point is further computed as the standard deviation of curvature between the point $x_{j}$ and its neighboring points $\mathcal{N}_{x_{j}}$ (with neighborhood size fixed as $k_{c}=10$, following prior work [23]), capturing local curvature fluctuations:
@@ 4.2 Patch-wise Spectral Trigger Injection
To overcome these limitations, we propose designing local patch-wise triggers in the spectral domain [14, 22]. Specially, a selected point cloud patch $S^{\prime}\in\{S^{\prime}_{1},S^{\prime}_{2},...,S^{\prime}_{m}\}$ is first transformed into an unweighted KNN graph $G=(V,A)$. Here $V\in\mathbb{R}^{k_{g}\times 3}$ represents the set of vertices, where each vertex corresponds to a point in the patch $S^{\prime}$. Following prior works [5, 14, 22], the adjacency matrix $A\in\mathbb{R}^{k_{g}\times k_{g}}$ is constructed by connecting each point to its $k_{p}=10$ nearest neighbors for graph construction, where $A_{ij}=1$ if points $i$ and $j$ are connected, otherwise $A_{ij}=0$.
@@
$\hat{S^{\prime}}=\phi_{GFT}(S^{\prime})=U^{T}S^{\prime},$ (4)
where $\hat{S^{\prime}}\in\mathbb{R}^{k_{g}\times 3}$. Correspondingly, the Inverse GFT (IGFT) could reconstruct the spatial feature from the spectral domain:
@@
Figure 3: Comparison of original and perturbed patches of airplane tail. The second row shows the spectral features of the original patch and spectral perturbations across different spectrum bands.
@@ 4.3 Optimization Strategy
As illustrated in Fig. 3, low-spectrum perturbations cause significant deformation to the patch of empennage, while high-spectrum perturbations introduce excessive outliers. Perturbations applied within intermediate spectral bands mitigate outliers while preserving the overall shape, presenting a key challenge for the optimal design of spectral triggers. To address this, we propose an optimizable spectral trigger $\xi$ that strikes a balance between attack effectiveness and stealthiness. The backdoor learning process follows an alternating optimization scheme, where the model parameters $\theta$ and the spectral trigger $\xi$ are updated iteratively, with one being optimized while the other remains frozen. Thus, each optimization step consists of two forward and backward propagations.
@@
5 Experiments and Results
5.1 Experiment Details
Attacks Setup. We compare our proposed SPBA method against four representative point cloud attack methods: PointBA-I [16], PointBA-O [16], IRBA [7], and IBAPC [5]. To reproduce PointBA-I, we implant a spherical cluster comprising 3% of the total points, with a fixed radius of 0.05, positioned at $(0.5,0.5,0.5)$ in the coordinate space. For PointBA-O, we apply a $10^{\circ}$ rotation along the Z-axis. For IRBA and IBAPC, we follow the default configurations specified in their original papers. For our proposed SPBA, $g$ and $k_{g}$ are both set as 32 for patch decomposition. And $m=16$ patches are selected for trigger injection. For the regularizer function in Eq. (10), we set $\lambda_{1}=1$, $\lambda_{2}=5$ and $\lambda_{3}=1$. However, for PointNet, we adjust $\lambda_{1}$ to 0.1 to improve stability.
For all attack methods, we set the poisoning rate $\rho=0.1$. The target class is chosen as Chair ($Y_{t}=8$) for ModelNe40 and Lamp ($Y_{t}=8$) for ShapeNetPart. We employ two Adam optimizers for backdoored model training and spectral trigger optimization, with learning rates of 0.001 and 0.01, respectively, and a weight decay of 1e-4. A cosine annealing learning rate scheduler is applied to both optimizers to facilitate stable convergence. Training is performed for 100 epochs with a batch size of 32. For all experiments, we use a workstation equipped with an NVIDIA RTX 4090 GPU.
Evaluation Metrics. The attack performance can be evaluated by Benign Accuracy (BA) and Attack Success Rate (ASR). BA is the accuracy of benign test samples that are correctly classified. ASR represents the proportion of poisoned test samples that are successfully misclassified into the target label. To quantify the imperceptibility of poisoned samples, we compute the Chamfer Distance (CD) between the poisoned and original samples. A successful backdoor attack should achieve a high ASR and lower CD while preserving BA as much as possible. For clarity, both BA and ASR are presented as percentages, and CD values in tables are scaled by a factor of 1,000.
@@ 5.1 Experiment Details
5.2 Attack Effectiveness
The results presented in Table 1 provide a comparative evaluation of various backdoor attack methods on ModelNet40. As shown in Table 1, our proposed SPBA achieves the lowest average CD value of 1.08 among all attacks, demonstrating its state-of-the-art stealthiness. What’s more, SPBA outperforms most backdoor attacks according to BA and ASR. In particular, compared to the sample-wise spectral attack IBAPC, SPBA achieves notable improvements of 0.47% in average BA and 2.78% in average ASR, highlighting the effectiveness of our patch-wise spectral trigger. Although PointBA-I achieves the highest average ASR of 98.43%, it comes at the cost of a significantly higher average CD of 4.08, making it much more detectable. Subsequent defense experiments further confirm that the trigger used in PointBA-I is especially susceptible to detection or removal by existing defense methods. Table 2 shows more experimental results on the ShapeNetpart dataset. SPBA achieves a similar performance, offering the best balance between attack effectiveness and stealthiness. These results further validate the superiority of SPBA as a more effective and imperceptible backdoor attack method compared to existing attacks.